OpenStudy (anonymous):

An airplane heading northwest at 500km/h encounters a wind of 120km/h from 25deg north of east. Determine the resultant ground velocity of the plane

OpenStudy (anonymous):

135 degres at 500km\h and 25 degres at 120km\h -use law of cos between both:135-25=110 resultant:r r^2=150^3+120^2-(2*500*120*cos110) r^2=305442.4172 r=552.67

OpenStudy (anonymous):

552.67km\h, if you draw a triangal, you can calculate the orientation

OpenStudy (anonymous):

where u get 135 deg from?